a sum of money compounded anuallly amount to rs.2340 in 8 years and 3185 in 10 years . find the annual rate of interest.
 
 

Grade:12th pass

1 Answers

Vikas TU
3 years ago
Dear student 
Dear student 
Compound interest = P(1+r/100)^n 
Applying the values we get 
2340(1+r/100)^2 = 3185 
(1+r/100)^2 = 3185/2340
(1+r/100)^2 = 49/36
1+r/100 = 7/6 
r/100 = 1/6 
r = 100/6 = 16.66%